# OpenHuEval: Evaluating Large Language Model on Hungarian Specifics

### v1

@@ Abstract
Haote Yang1thanks: Equal contribution., Xingjian Wei [^1] footnotemark: 1, Jiang Wu [^1] footnotemark: 1thanks: Project lead., Noémi Ligeti-Nagy211footnotemark: 1, Jiaxing Sun311footnotemark: 1, Yinfan Wang411footnotemark: 1, Zijian Győző Yang211footnotemark: 1, Junyuan Gao511footnotemark: 1, Jingchao Wang6, Bowen Jiang7, Shasha Wang1, Nanjun Yu6, Zihao Zhang8, Shixin Hong9, Hongwei Liu1, Wei Li1, Songyang Zhang1, Dahua Lin1,10, Lijun Wu1, Gábor Prószéky2, Conghui He1thanks: Corresponding author. 1Shanghai Artificial Intelligence Laboratory, 2HUN-REN Hungarian Research Centre for Linguistics, 3Wuhan University, 4Shanghai University, 5University of Chinese Academy of Sciences, 6East China Normal University, 7Peking University, 8Harbin Institute of Technology, 9Tsinghua University, 10Chinese University of Hong Kong Correspondence: heconghui@pjlab.org.cn
@@ 1 Introduction
2) Keeping up with the latest advances in LLM evaluation Significant progress has been made in LLM evaluation, with query sources shifting from manual or rule-based constructions to real-world internet questions Lin et al. (2024), enhancing practical relevance. Question formats evolved from multiple-choice to open-ended Q&A Wei et al. (2024), better reflecting actual usage. Evaluation methods transitioned from rule-based approaches to LLM-as-judge and subjective assessments, improving accuracy and objectivity Li et al. (2024). However, these advancements primarily apply to English datasets and not Hungarian. Thus, when creating OpenHuEval, we incorporated these principles and methodologies from English evaluations.
@@
- We established the framework for analyzing the thinking processes of the cutting-edge LRMs, revealing the intrinsic patterns and mechanisms of these models in non-English languages, with Hungarian as a representative.
@@
$\mathcal{L}$) Basic knowledge of the Hungarian language and Hungarian proverbs and sayings 1333
$\mathcal{H}$) Historical events and historical development of Hungary 324
@@
$\mathcal{F}$) Famous figures of Hungary 409
@@ 2.1 Hungarian-specific dimensions and OpenHuEval tasks
HuSimpleQA $\mathcal{L}$$\mathcal{H}$,$\mathcal{LCC}$,$\mathcal{EP}$,$\mathcal{GP}$$\mathcal{F}$,$\mathcal{PPL}$$\mathcal{BF}$ llm OE 1293
@@ 2.2 Hungarian WildBench
Task Introduction: The Hungarian WildBench (HuWildBench) task aims to evaluate the performance of LLMs in answering various questions arising from the everyday lives of Hungarians. All questions are sourced from Hungary’s well-known forum website [^1] https://www.gyakorikerdesek.hu/, which is similar to https://www.quora.com/ for English-speaking world. (hereinafter referred to as “g13k” for brevity). These questions cover a wide range of topics, including cultural customs, education, tourism, legal regulations, and business and finance, thus reflect real-life issues encountered by Hungarians. Examples of HuWildBench questions are shown in Figure 2 and Table 15. The queries in HuWildBench are user-generated content, therefore their linguistic expressions and question formats tend to be less formal than the structured and polished written language. This poses the realistic challenge for LLMs, as they must adapt to such informal and spontaneous language style. The construction of HuWildBench is detailed in Appendix B.
Metric and judge: We use the WB-Score Lin et al. (2024) as evaluation metric for HuWildBench, which is calculated in the following manner: We have simultaneously developped the customized checklist for each question during the process of creating each question, to assist the LLM judge in evaluating the responses. Examples of these checklists can be seen in Table 15. Following Lin et al. (2024), GPT-4o is used as the judge model, which evaluates the quality of each response based on the checklist and provides detailed strengths and weaknesses before assigning a score from 1 to 10. The definition of scores is shown in Table 6 and the judge prompt is detailed in Appendix B.5. Different from Lin et al. (2024), our final scores are calculated as the average of all test sample scores, with each score multiplied by 10.
@@ 2.3 Hungarian SimpleQA
Task Introduction: Hungarian SimpleQA (HuSimpleQA) is designed to assess the ability of LLMs to answer short, fact-seeking questions related to Hungary. Inspired by Wei et al. (2024) and He et al. (2024), we constructed HuSimpleQA based on Hungarian Wikipedia [^2] https://hu.wikipedia.org/, with the following key characteristics. Hungarian: The questions in HuSimpleQA are in Hungarian, and they focus on facts specifically related to Hungary. Diverse: The questions in HuSimpleQA cover the eight Hungary-specific dimensions proposed in §2.1. High-quality: The construction process of HuSimpleQA (in Appendix C) includes comprehensive and strict quality control procedures, ensuring the quality and accuracy of the questions. Static: Similar to SimpleQA, the answers to the questions in HuSimpleQA do not change over time, ensuring that the dataset remains evergreen. Easy-to-evaluate: The questions and answers in HuSimpleQA are short and concise, making them ideal for evaluation through LLMs. The examples of HuSimpleQA are shown in Figure 2 and Table 16. The construction of HuSimpleQA is detailed in Appendix C.
Metric and Judge: Following Wei et al. (2024), we use GPT-4o as a judge to categorize the responses of the LLM to HuSimpleQA into three classes: CORRECT, INCORRECT or NOT_ATTEMPTED. The judge prompt can be found in Appendix C.7, Figure 27.
@@ 2.4 Hungarian Proverb Reasoning
Task Introduction: Hungarian Proverb Reasoning (HuProverbRea), which consists of the collection of Hungarian proverbs, idioms, abbreviations, is a task that requires the LLM to understand and reason the meaning of Hungarian proverbs in a specific context. As shown by the examples in Figure 2, LLM is provided with a context in which a Hungarian proverb is used, accompanied by a question: “What does the speaker mean by the saying?”. Then, the LLM is tasked with discerning the speaker’s true intention, either by selecting the correct interpretation from two provided options (2CQ setting), or by directly articulating the speaker’s intended meaning (OE setting). The construction of HuProverbRea is detailed in Appendix D.
Metric and judge: For the 2CQ setting, we simply measure the correct ratio of candidate LLMs. For the OE setting, we adopt GPT-4o as judge to decide if the answer is acceptable. We provide the original proverb, its context and the English explanation of the proverb as references when judging OE responses. Detailed prompt templates are listed in Appendix D.
@@ 2.5 Hungarian Matching and Standard Filling-in-Blank
The constructions of HuMatchingFIB and HuStandardFIB are detailed in Appendix E.
Metric and Judge: In our inference prompts, we explicitly instruct the LLM to generate responses in a specified format. Additionally, we have established a set of well-defined rules to evaluate the correctness of the LLM’s answers for each blank. The detailed format requirements and judgment criteria can be found in the Appendix E.
@@ 3.3 Comparison with Existing Benchmarks
HuWildBench vs WildBench Lin et al. (2024): As shown in Figure 3(a), 70% LLMs experienced ranking changes. Detailed results can be found in Table  7.
@@ 4.2 Analysis on HuSimpleQA
The LRM’s responses to HuSimpleQA have been judged as CORRECT, INCORRECT, or NOT_ATTEMPTED. Then we used GPT-4o to split the thinking process into thoughts (see the prompt in Figure 37 and Figure 38). We evaluated the correctness of each thought (see the prompt in Figure 39), with examples provided Figure 40 and Figure 41 in Appendix F.
@@ 4.3 Analysis on HuMatchingFIB
We further tagged the reasoning segments according to the following four dimensions: (Dim1) correctness: Are the answers in this reasoning segment correct? (Dim2) complexity: In this reasoning segment, does the LRM simply assert the answer, or does it involve more complex reasoning? (Dim3) scope: Does this reasoning segment focuses on a single blank, modifies previous blanks, or addresses multiple blanks? (Dim4) language transfer: Does the LRM switch languages within this reasoning segment? The details of the tagging can be found in Appendix F.2.2. Examples of the tagging results can be found in Figure 47 and Figure 48.
@@
Explicit Translation Insertion (ETI): We observed that in some reasoning segments, when faced with the problem in Hungarian, the LLM first translates the key phrase of the original question into English and then proceeds with analysis and reasoning based on this translation. For example, “… Erőteljes #3# és a költői #4# gazdag használata jellemzi. This translates to "It is characterized by strong #3# and rich use of poetic #4#." …”. We refer to this phenomenon as Explicit Translation Insertion (ETI). Statistical analysis shows that ETI occurs in 5.49% of DeepSeek-R1’s reasoning segments, while for QwQ, the proportion is 16.77%. This indicates that QwQ is relatively weaker in handling non-English inputs, tending to translate first and then reason, which is consistent with previous research findings on cross-lingual Chain-of-Thought (CoT) in LLMs Sun et al. (2024).
@@ Acknowledgments
This research was supported by Shanghai Artificial Intelligence Laboratory.
@@ A.2 Language specific benchmarks for LLMs
Appendix B HuWildBench
B.1 Overall Construction Pipeline of HuWildBench
@@
(3) Deduplication: To ensure the diversity of questions, we performed deduplication on the Hungary-related questions within each secondary tag. The detailed process is outlined in Appendix B.2. After deduplication, the number of user questions in $Q_{dedup}$ was approximately 134K.
(4) Automatic high-quality question filtering: To ensure that only high-quality questions are extracted from the question pool, we designed a comparative-based high-quality question filtering strategy, as detailed in Appendix B.3. After filtering, the resulting set $Q_{high\_quality}$ contained around 2K questions.
@@
(6) Checklist construction: Based on WildBench Lin et al. (2024), we constructed a checklist for each question. The purpose of the checklist is to assist the LLM judger in evaluating the answers. Each item in the checklist queries a specific aspect of the answer to a question. An example of the checklist can be found in Table 15, and the detailed construction method is provided in Appendix B.4. To ensure the relevance of the checklist items to the questions, we hired a Hungarian native speaker to review the checklist for quality, filtering out non-compliant items and performing deduplication. The filtering criterion was whether the item was suitable as an evaluation dimension for the model’s response. To ensure the reliability of the LLM-as-judge, we filtered out user questions with fewer than 8 checklist items. The final set $Q_{checklist}$ contains 1154 questions. In the end, we obtained 1154 user questions along with their corresponding checklists.
B.2 Deduplication of similar questions
@@
B.3 Automatic high-quality question filtering
In order to automate the filtering of high-quality sample pots, we constructed the prompt that allows the GPT-4o to select the two best Hungarian questions out of the five based on the criteria of linguistic complexity, Hungarian relevance, common-sense accuracy, context-dependence, answer diversity, ambiguity, reasoning requirements, socio-ethical considerations, format diversity, and breadth of knowledge and outputs their indexes in JSON format to output their indexes, as shown in Figure 14. Specifically, we first set the criteria for high-quality questions in Prompt. Then we ask GPT-4o to compare the input questions based on the criteria. In order to mitigate the occurrence of some high-quality questions being eliminated prematurely (or vice versa) when all the questions in the same batch are of high quality, we follow the following 3 rules when filtering the high-quality questions: (1) filter 2 high-quality questions from 5 questions at a time, instead of filtering 1 high-quality question directly from 2 questions. (2) use the Swiss system mechanism instead of the knockout mechanism. In each screening round, each question can win in the current round as long as it ensures that it wins in two comparisons, and it will not be eliminated directly because of a failure in one comparison. (3) Our question screening strategy eliminates 65% of the questions in each round, in order to ensure that each secondary label has a sufficient number of high-quality questions. We conducted different elimination rounds for questions under different labels, and finally got about 2K questions. Finally, in order to validate the strategy of high-quality question screening, we manually checked about 200 5-option-2 results, and the pass rate was more than 80%, which proved the effectiveness of the present strategy. The final constructed HuWildBench is shown in Table 15.
B.4 Checklist construction

### v2

@@ Abstract
Haote Yang1thanks: Equal contribution., Xingjian Wei [^1] footnotemark: 1, Jiang Wu [^1] footnotemark: 1thanks: Project lead., Noémi Ligeti-Nagy211footnotemark: 1, Jiaxing Sun311footnotemark: 1, Yinfan Wang411footnotemark: 1, Zijian Győző Yang211footnotemark: 1, Junyuan Gao511footnotemark: 1, Jingchao Wang6, Bowen Jiang7, Shasha Wang1, Nanjun Yu6, Zihao Zhang8, Shixin Hong9, Hongwei Liu1, Wei Li1, Songyang Zhang1, Dahua Lin1,10, Lijun Wu1, Gábor Prószéky2, Conghui He1thanks: Corresponding author. heconghui@pjlab.org.cn 1Shanghai Artificial Intelligence Laboratory, 2HUN-REN Hungarian Research Centre for Linguistics, 3Wuhan University, 4Shanghai University, 5University of Chinese Academy of Sciences, 6East China Normal University, 7Peking University, 8Harbin Institute of Technology, 9Tsinghua University, 10Chinese University of Hong Kong
@@ 1 Introduction
2) Keeping up with the latest advances in LLM evaluation Significant progress has been made in LLM evaluation, with query sources shifting from manual or rule-based constructions to real-world internet questions Lin et al. (2024), enhancing practical relevance. Question formats evolved from multiple-choice to open-ended Q&A Wei et al. (2024), better reflecting actual usage. Evaluation methods transitioned from rule-based approaches to LLM-as-judge and subjective assessments, improving accuracy and objectivity Li et al. (2024a). However, these advancements primarily apply to English datasets and not Hungarian. Thus, when creating OpenHuEval, we incorporated these principles and methodologies from English evaluations.
@@
- We established the framework for analyzing the thinking processes of the cutting-edge LRMs, revealing the intrinsic patterns and mechanisms of these models in the Hungarian language and specifics, and providing a reference for research related to other non-English languages.
@@
Language($\mathcal{L}$) Basic knowledge of the Hungarian language and Hungarian proverbs and sayings 1333
History($\mathcal{H}$) Historical events and historical development of Hungary 324
@@
Figure($\mathcal{F}$) Famous figures of Hungary 409
@@ 2.1 Hungarian-specific dimensions and OpenHuEval tasks
HuSimpleQA $\mathcal{L}$$\mathcal{H}$$\mathcal{LCC}$$\mathcal{EP}$$\mathcal{GP}$$\mathcal{F}$$\mathcal{PPL}$$\mathcal{BF}$ llm OE 1293
@@ 2.2 Hungarian WildBench
Task Introduction: The Hungarian WildBench (HuWildBench) task aims to evaluate the performance of LLMs in answering various questions arising from the everyday lives of Hungarians. All questions are sourced from Hungary’s well-known forum website [^1] https://www.gyakorikerdesek.hu/, which is similar to https://www.quora.com/ for English-speaking world. (hereinafter referred to as “g13k” for brevity). These questions cover a wide range of topics, including cultural customs, education, tourism, legal regulations, and business and finance, thus reflect real-life issues encountered by Hungarians. Examples of HuWildBench questions are shown in Figure 2 and Table 21. The queries in HuWildBench are user-generated content, therefore their linguistic expressions and question formats tend to be less formal than the structured and polished written language. This poses the realistic challenge for LLMs, as they must adapt to such informal and spontaneous language style. The construction of HuWildBench is detailed in Appendix D.
Metric and judge: We use the WB-Score Lin et al. (2024) as evaluation metric for HuWildBench, which is calculated in the following manner: We have simultaneously developped the customized checklist for each question during the process of creating each question, to assist the LLM judge in evaluating the responses. Examples of these checklists can be seen in Table 21. Following Lin et al. (2024), GPT-4o is used as the judge model, which evaluates the quality of each response based on the checklist and provides detailed strengths and weaknesses before assigning a score from 1 to 10. The definition of scores is shown in Table 8 and the judge prompt is detailed in Appendix D.5. Different from Lin et al. (2024), our final scores are calculated as the average of all test sample scores, with each score multiplied by 10.
@@ 2.3 Hungarian SimpleQA
Task Introduction: Hungarian SimpleQA (HuSimpleQA) is designed to assess the ability of LLMs to answer short, fact-seeking questions related to Hungary. Inspired by Wei et al. (2024) and He et al. (2024), we constructed HuSimpleQA based on Hungarian Wikipedia [^2] https://hu.wikipedia.org/, with the following key characteristics. Hungarian: The questions in HuSimpleQA are in Hungarian, and they focus on facts specifically related to Hungary. Diverse: The questions in HuSimpleQA cover the eight Hungary-specific dimensions proposed in §2.1. High-quality: The construction process of HuSimpleQA (in Appendix E) includes comprehensive and strict quality control procedures, ensuring the quality and accuracy of the questions. Static: Similar to SimpleQA, the answers to the questions in HuSimpleQA do not change over time, ensuring that the dataset remains evergreen. Easy-to-evaluate: The questions and answers in HuSimpleQA are short and concise, making them ideal for evaluation through LLMs. The examples of HuSimpleQA are shown in Figure 2 and Table 22. The construction of HuSimpleQA is detailed in Appendix E.
Metric and Judge: Following Wei et al. (2024), we use GPT-4o as a judge to categorize the responses of the LLM to HuSimpleQA into three classes: CORRECT, INCORRECT or NOT_ATTEMPTED. The judge prompt can be found in Appendix E.7, Figure 27.
@@ 2.4 Hungarian Proverb Reasoning
Task Introduction: Hungarian Proverb Reasoning (HuProverbRea), which consists of the collection of Hungarian proverbs, idioms, abbreviations, is a task that requires the LLM to understand and reason the meaning of Hungarian proverbs in a specific context. As shown by the examples in Figure 2, LLM is provided with a context in which a Hungarian proverb is used, accompanied by a question: “What does the speaker mean by the saying?”. Then, the LLM is tasked with discerning the speaker’s true intention, either by selecting the correct interpretation from two provided options (2CQ setting), or by directly articulating the speaker’s intended meaning (OE setting). The construction of HuProverbRea is detailed in Appendix F.
Metric and judge: For the 2CQ setting, we simply measure the correct ratio of candidate LLMs. For the OE setting, we adopt GPT-4o as judge to decide if the answer is acceptable. We provide the original proverb, its context and the English explanation of the proverb as references when judging OE responses. Detailed prompt templates are listed in Appendix F.
@@ 2.5 Hungarian Matching and Standard Filling-in-Blank
The constructions of HuMatchingFIB and HuStandardFIB are detailed in Appendix G.
Metric and Judge: In our inference prompts, we explicitly instruct the LLM to generate responses in a specified format. Additionally, we have established a set of well-defined rules to evaluate the correctness of the LLM’s answers for each blank. The detailed format requirements and judgment criteria can be found in the Appendix G.
@@ 3.3 Comparison with Existing Benchmarks
HuWildBench vs WildBench Lin et al. (2024): As shown in Figure 3(a), 70% LLMs experienced ranking changes. Detailed results can be found in Table  9.
@@ 4.2 Analysis on HuSimpleQA
The LRM’s responses to HuSimpleQA have been judged as CORRECT, INCORRECT, or NOT_ATTEMPTED. Then we used GPT-4o to split the thinking process into thoughts (see the prompt in Figure 37 and Figure 38). We evaluated the correctness of each thought (see the prompt in Figure 39), with examples provided Figure 40 and Figure 41 in Appendix H.
@@ 4.3 Analysis on HuMatchingFIB
We further tagged the reasoning segments according to the following four dimensions: (Dim1) correctness: Are the answers in this reasoning segment correct? (Dim2) complexity: In this reasoning segment, does the LRM simply assert the answer, or does it involve more complex reasoning? (Dim3) scope: Does this reasoning segment focuses on a single blank, modifies previous blanks, or addresses multiple blanks? (Dim4) language transfer: Does the LRM switch languages within this reasoning segment? The details of the tagging can be found in Appendix H.2.2. Examples of the tagging results can be found in Figure 47 and Figure 48.
@@
Explicit Translation Insertion (ETI): We observed that in some reasoning segments, when faced with the problem in Hungarian, the LRM first translates the key phrase of the original question into English and then proceeds with analysis and reasoning based on this translation. For example, “… Erőteljes #3# és a költői #4# gazdag használata jellemzi. This translates to "It is characterized by strong #3# and rich use of poetic #4#." …”. We refer to this phenomenon as Explicit Translation Insertion (ETI). Statistical analysis shows that ETI occurs in 42.5% of DeepSeek-R1’s reasoning segments, while for QwQ, the proportion is 31.9%. This demonstrates that the LRM can adaptively translate input from the language in which it is relatively less proficient (such as Hungarian) into the more proficient language (such as English) during its thinking process. By scaling up the length of reasoning during the test stage, the LLM can partially overcome the cross-language performance discrepancies typical of traditional instruction-based LLM.
@@ Acknowledgments
This research was supported by National Key R&D Program of China (2022ZD0160201), and Shanghai Artificial Intelligence Laboratory.
@@ A.2 Language specific benchmarks for LLMs
Appendix B Hungarian-Specific Dimensions
OpenHuEval encompasses eight Hungarian-Specific Dimensions (HuSpecificDim), as shown in Table 2: Language ($\mathcal{L}$), History ($\mathcal{H}$), Life, Culture, and Customs ($\mathcal{LCC}$), Education and Profession ($\mathcal{EP}$), Geography and Place ($\mathcal{GP}$), Figure ($\mathcal{F}$), Politics, Policy, and Law ($\mathcal{PPL}$), and Business and Finance ($\mathcal{BF}$).
To select these Hungarian-Specific Dimensions, e first conducted extensive literature research to summarize the knowledge dimensions in existing cultural benchmarks, as shown in Table 6.
  Benchmark Dimension
CHARM Sun et al. (2024) History, Traditional Culture and Arts, Daily Life and Customs, Entertainment, Public Figures, Geography, Chinese Language
CUNIT Li et al. (2024b) Clothing, Food
BLEnD Myung et al. (2024) Food, Sports, Family, Education, Holidays/Celebrations/Leisure, Work-Life
CultureAtlas Fung et al. (2024) Culture, Holidays, Dining Etiquette, Education, Honorifics, etc.
GEOMLAMA Yin et al. (2022) Habits and Personal Choices, Cultures and Customs, Policies and Regulations, Geography
Table 6: The knowledge dimensions present in existing cultural benchmarks.
Based on above and by considering the knowledge needs of typical Hungarian users and the content on authoritative Hungarian websites, we established the Hu-specific dimensions for OpenHuEval. We believe these dimensions cover most user needs related to Hungarian specifics.
Appendix C Inference Detail
C.1 Default Maximum Output Length for LLM in Opencompass
For traditional instruction-based LLMs, we adopted OpenCompass’s default settings for the maximum output length, as shown in Table 7.
  Model Max out length
Qwen2.5-Instruct-7B 4096
Qwen2.5-Instruct-72B 4096
GPT4-o 2048
GPT-4o-mini 2048
Llama-3.1-Instruct-8B 1024
Llama-3.1-Instruct-70B 4096
Deepseek-V3 2048
Table 7: Default maximum output length for LLM in OpenCompass.
Appendix D HuWildBench
D.1 Overall Construction Pipeline of HuWildBench
@@
(3) Deduplication: To ensure the diversity of questions, we performed deduplication on the Hungary-related questions within each secondary tag. The detailed process is outlined in Appendix D.2. After deduplication, the number of user questions in $Q_{dedup}$ was approximately 134K.
(4) Automatic high-quality question filtering: To ensure that only high-quality questions are extracted from the question pool, we designed a comparative-based high-quality question filtering strategy, as detailed in Appendix D.3. After filtering, the resulting set $Q_{high\_quality}$ contained around 2K questions.
@@
(6) Checklist construction: Based on WildBench Lin et al. (2024), we constructed a checklist for each question. The purpose of the checklist is to assist the LLM judger in evaluating the answers. Each item in the checklist queries a specific aspect of the answer to a question. An example of the checklist can be found in Table 21, and the detailed construction method is provided in Appendix D.4. To ensure the relevance of the checklist items to the questions, we hired a Hungarian native speaker to review the checklist for quality, filtering out non-compliant items and performing deduplication. The filtering criterion was whether the item was suitable as an evaluation dimension for the model’s response. To ensure the reliability of the LLM-as-judge, we filtered out user questions with fewer than 8 checklist items. The final set $Q_{checklist}$ contains 1154 questions. In the end, we obtained 1154 user questions along with their corresponding checklists.
D.2 Deduplication of similar questions
@@
D.3 Automatic high-quality question filtering
In order to automate the filtering of high-quality sample pots, we constructed the prompt that allows the GPT-4o to select the two best Hungarian questions out of the five based on the criteria of linguistic complexity, Hungarian relevance, common-sense accuracy, context-dependence, answer diversity, ambiguity, reasoning requirements, socio-ethical considerations, format diversity, and breadth of knowledge and outputs their indexes in JSON format to output their indexes, as shown in Figure 14. Specifically, we first set the criteria for high-quality questions in Prompt. Then we ask GPT-4o to compare the input questions based on the criteria. In order to mitigate the occurrence of some high-quality questions being eliminated prematurely (or vice versa) when all the questions in the same batch are of high quality, we follow the following 3 rules when filtering the high-quality questions: (1) filter 2 high-quality questions from 5 questions at a time, instead of filtering 1 high-quality question directly from 2 questions. (2) use the Swiss system mechanism instead of the knockout mechanism. In each screening round, each question can win in the current round as long as it ensures that it wins in two comparisons, and it will not be eliminated directly because of a failure in one comparison. (3) Our question screening strategy eliminates 65% of the questions in each round, in order to ensure that each secondary label has a sufficient number of high-quality questions. We conducted different elimination rounds for questions under different labels, and finally got about 2K questions.
To ensure the reasonableness and robustness of the standard, we reviewed the results obtained using it. We selected 500 sets of questions filtered by GPT-4o according to this standard (where the model had to select two high-quality questions from each set of five). Human experts annotated these sets, and the overlap rate between the selections of the human experts and the model was 81% (a set was considered overlapping if the two questions chosen by the model matched those chosen by the experts). This demonstrates the effectiveness of the standard for selecting high-quality questions. The model’s recall rate for the questions selected by the human experts was 90%.
The final constructed HuWildBench is shown in Table 21.
D.4 Checklist construction